# Tweezer interferometry with NOON states

### v1

@@ Abstract
Atomic interferometers measure phase differences along paths with exceptional precision. Tweezer interferometry represents a novel approach for this measurement by guiding particles along predefined trajectories. This study explores the feasibility of using condensed bosons in tweezer interferometry. Unlike the factor $\sqrt{N}$ enhancement expected with classical ensembles, using NOON state interferometry can yield an enhancement by a factor of $N$. We propose a protocol for a tweezer-based NOON state interferometer that includes adiabatic splitting and merging of condensed bosons, followed by adiabatic branching for phase encoding. Our theoretical analysis focuses on the conditions necessary to achieve adiabaticity and avoid spontaneous symmetry breaking. Additionally, we demonstrate the feasibility of our proposed scheme and estimate the time required to perform these sweep processes.
@@ I Introduction
Interferometry [1] with matter waves [2, 3, 4] is a sensitive method for measuring spatial variations in a force field or differential acceleration. This technique involves splitting an atomic wave packet [8, 9, 10, 6, 7], allowing it to propagate along two different paths, and then recombining it to estimate the relative phase shift acquired along the paths. The specific state of the wave function after splitting influences the interferometer’s sensitivity. When the two parts of wave packet are entangled, the state is referred to as non-classical [11, 12, 13]. Interferometry with non-classical states enables improvement in relative phase sensitivity, which can approach the Heisenberg limit $1/N$, where $N$ is the number of particles. However, achieving this improvement in an experimentally feasible scenario is still an open challenge.
@@
In the majority of atom interferometers, splitting, mirroring and recombining occur in momentum space through the absorption of photons, with the wave packets undergoing ballistic motion in between these processes [14]. Recently, a guided atomic interferometry scheme based on optical tweezers [15] has been proposed [16, 17]. This method employs micro-optical traps to coherently split the wave function of atoms, guide them along selected trajectories, and then merge them to detect the interference pattern. The original proposal suggested preparing approximately ${N\sim 100}$ fermions simultaneously in different vibrational states of the tweezer, to then engage in the interferometric sequence in parallel. This approach offers a sensitivity enhancement by a factor of $\sqrt{N}$ compared to the single atom scenario.
@@
In this work, we explore the scenario of trapping bosons, specifically a Bose-Einstein condensate (BEC), as opposed to fermions. This naturally introduces the possibility for novel methods for interferometry [2, 18, 19, 20, 21]. We focus in this work on NOON state interferomerty, where the bosons are in a symmetric superposition of all atoms being in a BEC in one of the interferometer’s paths.
@@
We propose and analyze a complete guided interferometry scheme using NOON states. To generate these states, we exploit the tunability of the interaction strength [22, 23] and the confining potential. In the presence of attractive interactions, an adiabatic change from a single well to a double well generates a NOON state between the two wells. After separating the wells, and letting the wave packet in each of them to acquire a distinct phase, the challenge lies in recombining them to realistically estimate the phase difference. By ’realistic’ we mean that there should be no demand for accuracy in any experimental parameter that scales exponentially with $N$. Merely reversing the many-body splitting process does not meet this criterion. Instead, we introduce a recombination method that employs a two-mode adiabatic following, which translates the relative phase into the probability of finding all atoms at one of the interferometer’s exit ports. Importantly, this scheme requires that the dynamic range of control over system parameters scales only linearly with $N$. This aspect makes our proposal plausible for experimental realization.
@@
Outline.– In section II we summarize the basics of two mode interferometry. We then introduce our suggested protocol Fig. 1 in Section III. The Hamiltonian modeling of the system is presented in Section IV. Then in section V we provide results of simulations that motivate the further analysis and the suggested optimization procedure in Sections VI and VII. The feasibility of the protocol and its comparison to to other proposed schemes are discussed in Sections VIII and IX.
@@
Figure 1: The suggested interferometry protocol. The control parameters are $J(t)$ and $\Delta(t)$. Their time dependence is plotted in the bottom panel (black and blue lines, respectively). The number of particles $N$ and the interaction $U$ are kept constant. The critical value of $J$ (see text) is indicated by dashed line. We start with all the particles condensed in the lower symmetric orbital of the dimer. The stages are: (a) Decreasing $J$ leads to adiabatic splitting; (b) Phase encoding; (c) Increasing $J$ leads to partial or full merging; (d) Increasing $\Delta$ to a non-zero small value; (e) Decreasing $J$ again leads to adiabatic branching. In the proposed scheme (see text) $\dot{J}$ is adaptive. The 3 upper rows of panels display: The potential landscape [vertical axis is energy]; the Bloch sphere representation of the evolving state [sphere coordinates are $(S_{x},S_{y},S_{z})$]; and the schematic bifurcation diagram. Further explanation of the figure, and of the Bloch sphere representation, is provided throughout the following sections.
@@
Figure 2: Phase estimation. Red line indicates the expected $P(n)$ of our suggested interferometry protocol with idealized NOON states. For the sake of comparison, the black curve is the probability distribution that one would observe in standard interferometry. The phase can be estimated from the average $\left\langle n\right\rangle$, or better, by fitting to the expected $P(n)$. With our protocol, the estimation requires multiple runs to asses the probabilities $P(0)$ and $P(N)$.
@@ II Two-Mode Interferometry
Let us define our objective as measuring a potential difference, $V$, between two regions (“sites”). If we have a single particle in a superposition of being in the two sites, the accumulated phase difference during time $\tau_{\text{enc}}$, the so-called phase encoding stage [2], is $\varphi=V\tau_{\text{enc}}$. We use below units such that $\hbar=1$. When considering $N$ bosons in two sites, it is convenient to employ the basis states $|N{-}n,n\rangle$, indexed by the occupations of the two sites. Any initial state can be written as a superposition $|\psi\rangle=\sum_{n}c_{n}|N{-}n,n\rangle$. After phase encoding it evolves into
@@
Here, we omitted the $\phi_{n}$, because even in the presence of interaction ${\phi_{N}=\phi_{0}}$, assuming that the interaction is the same in both sites. In spite of its obvious benefits, NOON interferometry requires solving two challenges: preparation of the initial state, and performing the phase estimation.
@@
NorthPole $\equiv$ $|N,0\rangle\ =\ \frac{1}{\sqrt{N!}}(a_{1}^{{\dagger}})^{N}|0,0\rangle$ (3)
SouthPole $\equiv$ $|0,N\rangle\ =\ \frac{1}{\sqrt{N!}}(a_{2}^{{\dagger}})^{N}|0,0\rangle$ (4)
EastPole $\equiv$ $\frac{1}{\sqrt{N!2^{N}}}(a_{1}^{{\dagger}}+a_{2}^{{\dagger}})^{N}|0,0\rangle$ (5)
EvenCat $\equiv$ $\frac{1}{\sqrt{2}}\left[|N,0\rangle+|0,N\rangle\right]$ (6)
OddCat $\equiv$ $\frac{1}{\sqrt{2}}\left[|N,0\rangle-|0,N\rangle\right]$ (7)
It is important to realize that a $\pi/2$ rotation of NorthPole state, which is generated by $S_{y}$, yields an EastPole coherent state and not a cat state. The preparation of the latter can be performed using a non-linear splitter, as we explain below.
@@
In standard interferometry, one starts with a coherent state that is represented by a Gaussian-like distribution at the North pole of the Bloch sphere. A Ramsey $\pi/2$ pulse rotates it to the EastPole, where ${\varphi=0}$. During the time $t$ it rotates along the Equator to ${\varphi=Vt}$. A second Ramsey $\pi/2$ pulse follows. The obtained $P(n)$ is illustrated in Fig. 2, where the expectation value of $n$ is related to $\varphi$, namely,
$\left\langle n\right\rangle\ \equiv\ \sum_{n}nP(n)n\ =\ [1-\cos(\varphi)]\frac{N}{2}\ \ .$ (8)
@@
If one uses the same protocol, but replaces the coherent state by a NOON state, one gets a $P(n)$ whose line-shape is insensitive to $\varphi$. A more sophisticated protocol is required for both the NOON state preparation and for the phase estimation. We suggest such a protocol below and compare it to different alternatives in the concluding section.
III The suggested protocol
@@
A way to prepare an approximate NOON state has been suggested [6], and we shall refer to it as non-adiabatic splitting. The idea is to start with an East pole coherent state, and allow it to stretch dynamically such that after a certain time $P(n)$ is peaked at ${n{=}0}$ and ${n{=}N}$, and relatively negligible for other values of $n$.
@@
We shall refer to our procedure for preparing a NOON state as adiabatic splitting. Unlike the non-adiabatic version, it is not sensitive to timing issues. Furthermore, we shall see that it provides a higher quality NOON state, for which $P(n)$ is negligible for $n$ that is neither 00 nor $N$. For the phase estimation we use a biased version of the same process, which we call adiabatic branching.
@@
In this section, we briefly summarize our proposed interferometry protocol, depicted in Fig. 1. In the following sections, we will present a thorough analysis of it. The on-site interaction is denoted as $U$, with time units chosen such that $NU=1$. The time-dependent control parameters are the hopping frequency between the two sites, $J(t)$, and the energy bias between the two sites, $\Delta(t)$.
The protocol begins with the atoms in the EastPole coherent state, where all the bosons are condensed in the lowest orbital of the symmetric dimer. The goal of stage (a), during which $\Delta=0$, is to adiabatically transform this coherent state into an EvenCat superposition. This is the process that we call adiabtic splitting. We index the adiabatic levels $E_{\nu}$ such that the ground state is $\nu{=}1$. Then, the phase encoding stage (b) follows. This stage, by itself, is formally non-adiabatic and induces transitions between the quasi-degenerate symmetric state $\nu{=}1$ and the antisymmetric state $\nu{=}2$. Ideally, higher levels are not involved. Stages (c) and (d) merge the two wave packets. Stage (e), during which $\Delta\neq 0$, is what we called adiabatic branching. This means that if, at the end of stage (c), the state is an even state ($\nu=1$), it is mapped after stage (d) to $|N,0\rangle$, whereas if, due to phase accumulation, the state becomes an odd state ($\nu=2$), it is mapped to $|0,N\rangle$.
@@
More generally, for any given $\varphi$, this protocol yields a superposition of the $|N,0\rangle$ and $|0,N\rangle$ states, with the probabilities reflecting the relative phase. Therefore, $P(n)$ has peaks at $n=0$ and at $n=N$ (Fig. 2). To estimate $\varphi$, a single measurement is insufficient as it does not provide $P(n)$. One needs to perform $M\gg 1$ measurements to determine the relative height of the two peaks. An important property of our protocol is that at each run all atoms should be found at one and only one of the output ports. Any deviation from that rule signals a problem in the data point, which can then be discarded. Thus, our protocols offers a natural way to improve the quality of the results by post-selection.
@@
Figure 3: Energy landscape and wavepacket dynamics. The phasespace of the dimer system is the two-dimensional Bloch sphere whose embedding coordinates are ${(S_{x},S_{y},S_{z})}$. Upper panel: WKB energy contours ${\mathcal{H}_{\text{dimer}}=E_{\nu}}$ are plotted on the Bloch sphere. We use units of time such that $NU{=}1$. The coupling $J{=}0.4$ is larger than $J_{c}$, while $\Delta{=}0$ and $N=30$. Lower panel, right: The Husimi function of an approximate cat state that is produced using a non-adiabatic splitting protocol. The initial state (not displayed) is a condensate with ${N=30}$ particles, at the unstable (hyperbolic) East pole. We set ${J=0.5}$, hence the non-linearity (${u=2}$) is such that the seperatrix (black line) extends up to the North and South Poles. The time of the evolution is $\pi/\omega_{x}$, to maximize the splitting. Lower panel, left: Snapshot of the evolving Husimi function during a non-optimal adiabatic splitting of the same condensate at ${J(t)=0.1}$. The sweep rate is ${\dot{J}=1/8}$. The Bloch sphere is slightly rotated such that the North pole is fully visible (the hidden South pole supports the other half of the distribution). The white points represent a cloud that is evolved using classical equations of motion. The initial classical cloud is a circle of radius ${n_{0}=1-(4N)^{-1}}$. The outcome of this simulation when ${J(t)=0}$ is approximately a NOON state with probability ${P(n)\approx 0.23}$ for getting either $n{=}0$ or $n{=}N$. For an ideal protocol one aims at getting ${P(N)\approx P(0)\approx 0.5}$.

### v2

@@ Abstract
Atomic interferometers measure phase differences along paths with exceptional precision. Tweezer interferometry represents a novel approach for this measurement by guiding particles along predefined trajectories. This study explores the feasibility of using condensed bosons in tweezer interferometry. Unlike the factor $\sqrt{N}$ enhancement expected with classical ensembles, using NOON state interferometry can yield an enhancement by a factor of $N$. We consider a protocol for a tweezer-based NOON state interferometer that includes adiabatic splitting and merging of condensed bosons, followed by adiabatic branching for phase encoding. Our theoretical analysis focuses on the conditions necessary to achieve adiabaticity and avoid spontaneous symmetry breaking. Additionally, we demonstrate the feasibility of the proposed scheme and estimate the time required to perform these sweep processes.
@@ I Introduction
Interferometry with matter waves [1, 2, 3] is a sensitive method for measuring spatial variations in a force field or differential acceleration. This technique involves splitting an atomic wave packet [7, 8, 9, 5, 6], allowing it to propagate along two different paths, and then recombining it to estimate the relative phase shift acquired along the paths. The specific state of the wave function after splitting influences the interferometer’s sensitivity. When the two parts of the wave packet are entangled, the state is referred to as non-classical [10, 11, 12]. Interferometry with non-classical states enables improvement in relative phase sensitivity, which can approach the Heisenberg limit $1/N$, where $N$ is the number of particles, see [13]. However, achieving this improvement in an experimentally feasible scenario is still an open challenge, mainly addressed so far in photonic systems [14, 15, 16].
@@
In the majority of atom interferometers, splitting, mirroring and recombining occur in momentum space through the absorption of photons, with the wave packets undergoing ballistic motion in between these processes [2]. Recently, a guided atomic interferometry scheme based on optical tweezers has been proposed [17, 18, 19]. This method employs micro-optical traps to coherently split the wave function of atoms, guide them along selected trajectories, and then merge them to detect the interference pattern. The original proposal [17] suggested preparing approximately ${N\sim 100}$ fermions simultaneously in different vibrational states of the tweezer, to then engage in the interferometric sequence in parallel. This approach offers a sensitivity enhancement by a factor of $\sqrt{N}$ compared to the single atom scenario.
@@
In this work, we explore the scenario of trapping bosons, specifically a Bose-Einstein Condensate (BEC), as opposed to fermions. This naturally introduces the possibility for novel methods for interferometry [1, 20, 21, 22, 23]. We focus in this work on NOON state interferomerty, where the bosons are in a symmetric superposition of all atoms being in a BEC in only one of the interferometer’s paths.
@@
We analyze a complete guided interferometry scheme using NOON states [24]. To generate these states, one exploits the tunability of the interaction strength [25, 26] and the confining potential. In the presence of attractive interactions, an adiabatic change from a single well to a double well generates a NOON state between the two wells. After separating the wells, and letting the wave packet in each of them to acquire a distinct phase, the challenge lies in recombining them to realistically estimate the phase difference. By ’realistic’ we mean that there should be no demand for accuracy in any experimental parameter that scales exponentially with $N$. Merely reversing the many-body splitting process does not meet this criterion. Instead, we incorporate a recombination method that employs a two-mode adiabatic following, which translates the relative phase into the probability of finding all atoms at one of the interferometer’s exit ports. Importantly, this scheme requires that the dynamic range of control over system parameters scales only linearly with $N$. This aspect makes our proposal plausible for experimental realization.
@@
Outline.– In section II we summarize the basics of two mode interferometry. We then introduce the protocol of Fig. 1 in Section III. The Hamiltonian modeling of the system is presented in Section IV. Then in section V we provide results of simulations that motivate the further analysis and the suggested optimization procedure in Section VI and Appendix B. The feasibility of the protocol and its comparison to to other proposed schemes are discussed in Sections VII and VIII.
@@
Figure 1: The interferometry protocol. The control parameters are $J(t)$ and $\Delta(t)$. Their time dependence is plotted in the bottom panel (black and blue lines, respectively). The number of particles $N$ and the interaction $U$ are kept constant. The critical value of $J$ (see text) is indicated by dashed line. We start with all the particles condensed in the lower symmetric orbital of the dimer. The stages are: (a) Decreasing $J$ leads to adiabatic splitting; (b) Phase encoding; (c) Increasing $J$ leads to partial or full merging; (d) Increasing $\Delta$ to a non-zero small value; (e) Decreasing $J$ again leads to adiabatic branching. In the proposed scheme (see text) $\dot{J}$ is adaptive. The 3 upper rows of panels display: The potential landscape, where the vertical axis is energy. Note that the horizontal axis is $S_{z}$, the occupation imbalance, and not real space; the Bloch sphere representation of the evolving state [sphere coordinates are $(S_{x},S_{y},S_{z})$]; and the schematic bifurcation diagram. Further explanation of the figure, and of the Bloch sphere representation, is provided throughout the following sections.
@@
Figure 2: Phase estimation. Red line indicates the expected $P(n)$ of the suggested interferometry protocol with idealized NOON states. For the sake of comparison, the black curve is the probability distribution that one would observe in standard interferometry. The phase can be estimated from the average $\left\langle n\right\rangle$, or better, by fitting to the expected $P(n)$. With our protocol, the estimation requires multiple runs to asses the probabilities $P(0)$ and $P(N)$, see Sec. (III) for details.
@@ II Two-Mode Interferometry
We define our objectiv as measuring a potential difference, $V$, between two regions (“sites”). If we have a single particle in a superposition of being in the two sites, the accumulated phase difference during time $\tau_{\text{enc}}$, the so-called phase encoding stage [1], is $\varphi=V\tau_{\text{enc}}$, where we employ units such that $\hbar=1$. When considering $N$ bosons in two sites, it is convenient to employ the basis states $|N{-}n,n\rangle$, indexed by the ground state occupation of the two sites. Any initial state can be written as a superposition $|\psi\rangle=\sum_{n}c_{n}|N{-}n,n\rangle$. After phase encoding it evolves into
@@
Here, we omitted the $\phi_{n}$, because even in the presence of interaction ${\phi_{N}=\phi_{0}}$, assuming that the interaction is the same in both sites. In spite of its obvious benefits, NOON interferometry requires overcoming two challenges: preparation of the initial state, and performing the phase estimation.
@@
$\left|Z\right\rangle$ $=$ $\text{NorthPole}\equiv|N,0\rangle\ =\ \frac{1}{\sqrt{N!}}(a_{1}^{{\dagger}})^{N}|0,0\rangle$ (3)
$\left|\bar{Z}\right\rangle$ $=$ $\text{SouthPole}\equiv|0,N\rangle\ =\ \frac{1}{\sqrt{N!}}(a_{2}^{{\dagger}})^{N}|0,0\rangle$ (4)
$\left|X\right\rangle$$\equiv$ $\frac{1}{\sqrt{N!2^{N}}}(a_{1}^{{\dagger}}+a_{2}^{{\dagger}})^{N}|0,0\rangle$ (5)
$\left|\bar{X}\right\rangle$$\equiv$ $\frac{1}{\sqrt{N!2^{N}}}(a_{1}^{{\dagger}}-a_{2}^{{\dagger}})^{N}|0,0\rangle$ (6)
[…]
@@
In standard interferometry, one starts with a coherent state that is represented by a Gaussian-like distribution at the North pole of the Bloch sphere. A Ramsey $\pi/2$ pulse rotates it to $\left|X\right\rangle$, where ${\varphi=0}$. During the time $t$ it rotates along the Equator to ${\varphi=Vt}$. A second Ramsey $\pi/2$ pulse follows. The obtained $P(n)$ is illustrated in Fig. 2, where the expectation value of $n$ is related to $\varphi$, namely,
$\left\langle n\right\rangle\ \equiv\ \sum_{n}nP(n)\ =\ [1-\cos(\varphi)]\frac{N}{2}\ \ .$ (9)
@@
If one uses the same protocol, but replaces the coherent state by a NOON state, one gets a $P(n)$ whose line-shape is insensitive to $\varphi$. A more sophisticated protocol is required for both the NOON state preparation and for the phase estimation. We describe such protocol in the next section, and compare it to different alternatives in the concluding section.
III The protocol
@@
A way to prepare an approximate NOON state has been suggested in Ref. [5], and we shall refer to it as non-adiabatic splitting. The idea is to start with an $\left|X\right\rangle$ coherent state, and allow it to stretch dynamically such that after a certain time $P(n)$ is peaked at ${n{=}0}$ and ${n{=}N}$, and relatively negligible for other values of $n$. The advantage of such protocol is that the evolution is generated by a time-independent Hamiltonian: it is a rapid quench process. The disadvantage is the need to have a precise control over the time duration of the quench, and the relatively low fidelity of the generated NOON state. The latter point is further discussed in the concluding section.
@@
The optional procedure for preparing a NOON state that we consider below is adiabatic splitting, as in [24]. Unlike the non-adiabatic quench, it is not sensitive to timing issues. Furthermore, we shall see that it provides a higher fidelity NOON state, for which $P(n)$ is negligible for $n$ that is neither 00 nor $N$. For the phase estimation we use the biased version of the same process, which we call adiabatic branching.
@@
In this section, we briefly summarize the interferometry protocol, depicted in Fig. 1. In the following sections, we will present a thorough analysis of it. The on-site interaction is denoted as $U$, with time units chosen such that $NU=1$. The time-dependent control parameters are the hopping (or tunneling) frequency between the two sites, $J(t)$, and the energy bias between the two sites, $\Delta(t)$. We index the adiabatic levels $E_{\nu}$ such that the ground state is $\nu{=}1$.
III.1 Stage (a) - adiabatic splitting
The protocol begins with the atoms in the $\left|X\right\rangle$ state as defined in Eq. (5). This is a coherent state, where all the bosons are condensed in the lowest orbital of the symmetric dimer. In the Bloch-Sphere language that we further elaborate in the next section, this state is represented by a Gaussian-like Hussimi function that is located at the bottom of a single-well potential. This potential (on the Bloch Sphere) should not be confused with the actual potential in space. Namely, we assume that the atoms are trapped at any time by a potential created by two optical tweezers that generate two potential minima in real space. In this work we model only the case of two separate minima. The control over the tunnel coupling $J$ can be achieved by changing the distance between the minima or the depth of each of the potentials.
[…]
III.2 Stage (b) - phase encoding
During the phase encoding stage the external bias leads to the accumulation of phase difference $N\varphi$ between the two sites. Thus, at the end of the phase encoding stage the state of the system is
$\left|\varphi\right\rangle\ \ =\ \ \frac{1}{\sqrt{2}}\left[\left|Z\right\rangle+e^{iN\varphi}\left|\bar{Z}\right\rangle\right]$ (10)
[…]
III.3 Stage (c) - merging
During the merging stage $J$ is increased, and the double well on the Bloch Sphere becomes again a single well (no self-trapping). In the absence of phase accumulation the adiabatic evolution leads back to the ground state $\left|X\right\rangle$, meaning that all the $N$ particles are condensed in the lower orbital. If we have e.g. phase accumulation ${N\varphi=\pi}$, it means that the system has evolved into the second adiabatic state ${\nu=2}$, and therefore after merging, we have $N{-}1$ particles in the lower orbital, and one particle in the excited orbital. More generally, due to phase accumulation the system ends up in a superposition of the first and second adiabatic levels. The first adiabatic level ($\nu=1$) is the ground state (condensation in the lowest orbital), while the second adiabatic level ($\nu=2$) is a one-particle excitation.
III.4 Stage (d) - Setting bias
Stage (d) looks rather innocent. A very small bias is applied, meaning that there is a potential difference $\Delta$ between the two sites. Using Bloch Sphere terminology it means that the single well minimum is no longer at ${S_{x}=N/2}$, but slightly shifted. This small bias has no significant effect as long as $J$ is kept large. It becomes important only as $J$ is decreased.
III.5 Stage (e) - adiabatic branching
The branching stage is formally similar to the splitting stage, but here the bias is non-zero ($\Delta\neq 0$). If, at the end of stage (c), the state is an even state ($\nu=1$), it is mapped after stage (e) to $|N,0\rangle$, whereas if, due to phase accumulation, the state at the end of stage (c) becomes an odd state ($\nu=2$), it is mapped to $|0,N\rangle$.
More generally, for any given $\varphi$, this protocol yields a superposition of the $|N,0\rangle$ and $|0,N\rangle$ states, with the probabilities reflecting the relative phase. Therefore, $P(n)$ has peaks at $n=0$ and at $n=N$, see Fig. 2.
III.6 Phase estimation
To estimate $\varphi$, a single measurement is insufficient as it does not provide $P(n)$. One needs to perform $M\gg 1$ measurements to determine the relative height $P(N)/P(0)$ of the two peaks in Fig. 2. This does not contradict the well known statement [13] that NOON state interferometry achieves the Heisenberg limit. The sensitivity scales as $1/N$ because the accumulated phase in the phase encoding stage is $N\varphi$. But in some sense there is also an $M$ dependent prefactor that determines the accuracy.
@@ III.6 Phase estimation
An important property of our protocol is that at each run all atoms should be found at one and only one of the output ports. Any deviation from that rule signals a problem in the data point, which can then be discarded. Thus, the protocol implies a natural way to improve the quality of the results by post-selection.
@@
Figure 3: Energy landscape and wavepacket dynamics. The phasespace of the dimer system is the two-dimensional Bloch sphere whose embedding coordinates are ${(S_{x},S_{y},S_{z})}$. Upper panel: WKB energy contours ${\mathcal{H}_{\text{dimer}}=E_{\nu}}$ are plotted on the Bloch sphere. We use units of time such that $NU{=}1$. The coupling $J{=}0.4$ is larger than $J_{c}$, while $\Delta{=}0$ and $N=30$. Lower panels: Husimi functions – probability to find the system in an ${(S_{x},S_{y},S_{z})}$ coherent state is indicated by color. Right panel: The Husimi function of an approximate cat state that is produced using a non-adiabatic splitting protocol. The initial state (not displayed) is a condensate $\left|X\right\rangle$ with ${N=30}$ particles, located at an unstable (hyperbolic) point of the Bloch sphere. We set ${J=0.5}$, hence the non-linearity (${u=2}$) is such that the separatrix (black line) extends up to the North and South Poles. The time of the evolution is $\pi/\omega_{x}$, to maximize the splitting. Left panel: Snapshot of the evolving Husimi function during a non-optimal adiabatic splitting of the same condensate at ${J(t)=0.1}$. The sweep rate is ${\dot{J}=1/8}$. The Bloch sphere is slightly rotated such that the North pole is fully visible (the hidden South pole supports the other half of the distribution). The white points represent a cloud that is evolved using classical equations of motion. The initial classical cloud is a circle of radius ${n_{0}=1-(4N)^{-1}}$. The outcome of this simulation when ${J(t)=0}$ is approximately a NOON state with probability ${P(n)\approx 0.23}$ for getting either $n{=}0$ or $n{=}N$. For an ideal protocol one aims at getting ${P(N)\approx P(0)\approx 0.5}$.